# Neural Networks and Friction: Slide, Hold, Learn

### v3

@@ Abstract
In this study, it is demonstrated that Recurrent Neural Networks (RNNs), specifically those utilizing Gated Recurrent Unit (GRU) architecture, possess the capability to learn the complex dynamics of rate-and-state friction laws from synthetic data. The data employed for training the network is generated through the application of traditional rate-and-state friction equations coupled with the aging law for state evolution. A novel aspect of our approach is the formulation of a loss function that explicitly accounts for initial conditions, the direct effect, and the evolution of state variables during training. It is found that the RNN, with its GRU architecture, effectively learns to predict changes in the friction coefficient resulting from velocity jumps, thereby showcasing the potential of machine learning models in understanding and simulating the physics of frictional processes.
@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3, 4, 5]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [6].
Friction modeling [7, 8] aims to capture essential features observed in experimental settings [9], notably the direct effect and healing [2]. The direct effect refers to the immediate change in frictional resistance to a sudden change in sliding velocity. This phenomenon can be described by the empirical equation [6]:
$\mu=\mu_{0}+a\log\left({V\over V_{ref}}\right)\,,$ (1)
where $\mu$ is the friction coefficient, $\mu_{0}$ is the friction coefficient at a reference velocity $V_{ref}$, $V$ is the sliding velocity, and $a$ is a positive constant representing the direct effect’s sensitivity to velocity changes [10].
Healing, also referred to as state evolution effect, illustrates the time-dependent increase in frictional resistance during stationary contact, attributed to the contact area’s growth at asperity contacts or chemical bonding at the interface [2]. This process is often modeled by the evolution of an internal variable $\theta$ (termed “state”), with the simplest form being:
$\dot{\theta}=1-\frac{\theta V}{D_{c}}\,,$ (2)
where $D_{c}$ is a critical slip distance over which the state variable evolves significantly [11]. The state is interpreted as a measure of effective age of contacts between the asperities of the opposed surfaces [10, 12], and it therefore has units of time. Equation 2 is termed “aging law”.
Numerous models have been proposed to describe frictional behavior, from classical laws like Amontons-Coulomb friction [2], which simplifies friction as a constant proportion of the normal load, to more complex models incorporating velocity and time-dependent effects. However, these models often fall short in capturing the full range of observed behaviors [6, 13].
Rate-and-state friction (RSF) laws stand out by explicitly addressing the direct effect and healing, thereby providing a more comprehensive framework for understanding the seismic cycle [14]. A customary expression of this law is [6]:
@@
where $b$ is a constant that, along with $a$, determines the frictional response to changes in sliding velocity and state evolution [15]. Despite their advantages, RSF laws are not without limitations. Identifying the parameters $a$, $b$, and $D_{c}$ from experimental data poses significant challenges due to the complex interplay between different physical processes at the frictional interface [6].
Criticism has been levied against the use of the “state” variable in RSF laws for its empirical nature and lack of a clear physical basis, leading to calls for models that include more explicit physical mechanisms [16, 17]. To begin, the parameters that appear in eq. 3 are assumed to be constants, and they may not [18, 19, 20, 21, 22]. Furthermore, the need to regularize these laws to address numerical instabilities and ensure the physical realism of simulations has been highlighted [23, 24]. Recently, laboratory [25] and virtual experiments [26, 27] have shown the limitations of RSF laws even in highly controlled interface conditions.
@@
Given these challenges and the limitations of existing models, there is a clear need to explore new methodologies for understanding friction dynamics. This necessity is particularly acute for modeling phenomena that span multiple scales of time and space [28]. Herein lies the potential of leveraging advanced computational approaches, such as Recurrent Neural Networks (RNNs) [29], to offer new insights and overcome the limitations of traditional friction laws.
Neural networks have revolutionized the field of machine learning, offering powerful tools for modeling complex, non-linear relationships in data across a myriad of disciplines [30]. At their core, neural networks are composed of layers of interconnected nodes or ”neurons,” each capable of performing simple computations. Through the process of training, these networks adjust their internal parameters to minimize the difference between their output and the desired outcome, effectively learning to map inputs to outputs [31].
@@
Two notable advancements in RNN architecture are the Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) networks. Both GRU and LSTM address a fundamental limitation of basic RNNs known as the vanishing gradient problem, which makes it difficult for RNNs to learn dependencies between events that occur at long intervals in the input data [33, 31]. GRUs simplify the LSTM architecture while retaining its ability to capture long-term dependencies, making it both efficient and powerful for a wide range of tasks [32].
@@
This success in modeling path-dependent phenomena in materials science provides a strong foundation for the belief that similar methodologies can be applied with success to the study of friction. Friction, much like material plasticity, is inherently dependent on history, in its case of contact and slip between surfaces. The capability of RNNs to capture and learn from sequential, history-dependent data makes them promising candidates for advancing our understanding and modeling of friction dynamics.
@@
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular, eq. 2. This approach is essentially different from the one used recently by Ishiyama and collaborators [45], who used machine learning (random forest [46]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[17, 47, 48]. It also differs from the use made by Rouet-Leduc et al [49] and Hulbert et al [50], among others [51, 52], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. To the best of our knowledge, the closest work to ours may be the one by Wang and Sun[53], which focused on interfacial traction-opening laws, and the recent one by Cravero and Bhattacharya [54], who studied interface adhesion. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without explicitly defining internal variables [43], since the state is not a physical quantity that can be measured experimentally, cf. work by Kumar and collaborators [55, 56]. In reality, the complexities of interface physics [57] are such that probably no simplified internal variable can be defined as long as one aims at capturing any modestly ambitious degree of complexity. We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated by the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
@@
The structure of the paper is as follows: the methods section (1) introduces the GRU architecture, (2) describes the generation of synthetic data for training and verification, and (3) explains the choice of training hyperparameters, including the physics-informed loss function. The next section describes the outcome of the training process, which details performance evaluation, and then assesses the model’s predictive accuracy on independent datasets. The discussion section analyzes the findings and their implications. The final remarks section summarizes the study’s contributions and suggests future research directions.
@@ II.1 Neural networks for sequences: gated recurrent units
RNNs are engineered to handle sequential data thanks to their ability to maintain a memory of previous inputs by feeding the output of a neuron back onto itself [29]. This recursive nature allows making predictions based on the information accumulated from prior inputs. However, standard RNNs often struggle with long-term dependencies due to the vanishing gradient problem [31], where gradients become too small or too large for the network to learn effectively over many time steps.
Gated Recurrent Units (GRUs) are an evolution of the traditional RNN architecture, designed to mitigate the vanishing gradient problem and improve the network’s ability to learn from long sequences. GRUs introduce gating units that control the flow of information. These gates “decide” which information is relevant to keep or discard as the sequence progresses, enabling the network to maintain a longer memory.
The GRU mechanism is articulated mathematically through [32]:
  • Update gates:
    $z_{t}=\sigma(W_{z}x_{t}+U_{z}h_{t-1}+\xi_{z})\,.$ (4)
  • Reset gates:
    $r_{t}=\sigma(W_{r}x_{t}+U_{r}h_{t-1}+\xi_{r})\,.$ (5)
  • Candidate states:
    $\tilde{h}_{t}=\tanh(Wx_{t}+U(r_{t}\odot h{t-1})+\xi)\,.$ (6)
  • Final state updates:
    $h_{t}=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}\,.$ (7)
In these equations, $x_{t}$ represents the input at time step $t$; $h_{t}$ is the hidden state at time $t$; $\sigma$ denotes the sigmoid activation function (to ensure the gate activations are constrained between 0 and 1); $\tanh$ is the hyperbolic tangent function (introducing non-linearity); $W$, $U$, and $\xi$ are the weights and biases associated with the input and hidden states, respectively; and $\odot$ signifies element-wise multiplication.
By merging the forget and input gates into a single update gate, GRUs simplify the LSTM architecture, reducing the model’s complexity and computational demands while maintaining robust performance.
@@ II.2 Synthetic data generation
As mentioned earlier, it is assumed that the state $\theta$ will not be needed by the neural network that is to replace the traditional RSF law. This means that there will not be vectors of values of $\theta$ being used during training and that the network will not , but it does not imply that eq. 2 is not being used during the datasets’ generation. Let us reiterate that we assume a hypothetical interface that is perfectly described by eq. 3 in tandem with eq. 7, but this fact would be hidden to us if we were to analyze experimental data coming out of that interface.
So, to generate the dataset for training and verifying our model, we start by defining a traditional RSF model with specific parameter values, grounded in the fundamental equations presented earlier (eq. 3), with parameter values set as follows:
@@
Using these values, we simulate 15000 random sequences to reflect diverse frictional sliding scenarios. Each sequence is designed to simulate a total sliding distance of $30D_{c}$ and is completed within 0.013 seconds. The number of velocity jumps within each sequence is randomly determined, drawing from a uniform distribution with values 2, 3, 4, 5. It is assumed that each sequence initiates from a non-zero velocity, with the initial state of the interface corresponding to the steady state for that velocity, as defined by the RSF laws: $\theta_{ss}=D_{c}/V(t=0)$.
Velocity jumps within these sequences occur at random intervals, determined by sampling times from a uniform distribution across the interval 0, 1, with 0 marking the start and 1 the end of the sliding protocol. Each sequence incorporates a hold period where the velocity drops to zero, simulating intervals where sliding stops until the subsequent jump. To avoid errors associated to dividing by zero or trying to compute the logarithm of zero, a tiny yet finite velocity is assigned to the hold for computations ($2\cdot 10^{-9}$ m/s).
@@
Having the velocity protocol, we proceed to calculate the evolution of the state variable $\theta$ using eq. 2 and initial conditions $V(t=0)=V_{0}$ and $\theta(t=0)=\theta_{ss}$. With the velocity protocol and the evolving state variable at hand, we can then compute the friction coefficient’s evolution over time, as described by the RSF model (eq. 3).
The dataset comprises pairs of “features” and “targets”. Features represent the velocity protocol sampled at 100 equally-spaced instants during the duration of each virtual experiment, while targets correspond to variations of the friction coefficient sequence at these same instants. What is the reason for working with variations and not with the actual values of friction coefficient? There is no way to predict the initial values of friction based on these datasets, they have to be provided as initial conditions over which the variations are later superimposed.
For training, we normalize using the following characteristic values: $\delta\mu_{ch}=0.01$ for the friction coefficient change and $V_{ch}=0.3$ m/s for the velocities. $20\%$ of the sequences are put aside to test the network once it is trained.
@@ II.3 Network details
Mathematica’s [58] neural networks package has been used.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a single GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation.
Figure 2: Scheme of main part of the network

### v4

@@ Abstract
In this study, it is demonstrated that Recurrent Neural Networks (RNNs), specifically those utilizing Gated Recurrent Unit (GRU) architecture, possess the capability to learn the complex dynamics of rate-and-state friction laws from synthetic data. The data employed for training the network is generated through the application of traditional rate-and-state friction equations coupled with either the aging law or the for state evolution. A novel aspect of our approach is the formulation of a loss function that explicitly accounts for the direct effect by means of automatic differentiation. It is found that the RNN, with its GRU architecture, effectively learns to predict changes in the friction coefficient resulting from velocity jumps (with and without noise in the target data), thereby showcasing the potential of machine learning models in capturing and simulating the physics of frictional processes.
@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3, 4, 5, 6]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [7].
Numerous models have been proposed to describe frictional behavior, from classical laws like Amontons-Coulomb friction [2], which simplifies friction as a constant proportion of the normal load, to much more complex ones. However, these models often fall short in capturing the full range of observed behaviors [7, 8].
Rate-and-state friction (RSF) laws stand out by explicitly addressing the velocity and time-dependent effects, thereby providing a more comprehensive framework for understanding the seismic cycle [9]. A customary expression of this law is [7]:
@@
where $\mu$ is the friction coefficient, $\mu_{0}$ is the friction coefficient at a reference velocity $V_{ref}$, $V$ is the sliding velocity, and $a$ is a positive constant representing the direct effect’s sensitivity to velocity changes [10], $b$ is a constant that, along with $a$, determines the frictional response to changes in sliding velocity and state evolution [11].
Friction modeling [12, 13] aims to capture essential features observed in experimental settings [14], notably the direct effect and healing [2].
Healing, also referred to as state evolution effect, illustrates the time-dependent increase in frictional resistance during stationary contact, attributed to the contact area’s growth at asperity contacts or chemical bonding at the interface [2]. This process is often modeled by the evolution of an internal variable $\theta$ (termed “state”), with the simplests form being the “aging law” and the “slip law”:
$\dot{\theta}={d\theta\over dt}$ $=1-\frac{\theta V}{D_{c}}\,,$ (2a)
$\dot{\theta}={d\theta\over dt}$ $=-\frac{\theta V}{D_{c}}\log\left(\frac{\theta V}{D_{c}}\right)\,,$ (2b)
where $t$ is time and $D_{c}$ is a critical slip distance over which the state variable evolves significantly [15]. In the aging law, the state is interpreted as a measure of effective age of contacts between the asperities of the opposed surfaces [10, 16], a logical position given that contact area between two rough surfaces is also to grow logarithmically over time during periods of rest or very slow sliding [5, 17]. Notwithstanding this intuitive inference, the slip law offers a picture that may be better backed by experiments [18, 19], in which the sensitivity to sliding does not ever disappear, not even at tiny sliding rates [19].
Despite their advantages, RSF laws are not without limitations. Identifying the parameters $a$, $b$, and $D_{c}$ from experimental data poses significant challenges due to the complex interplay between different physical processes at the frictional interface [7].
Criticism has been levied against the use of the “state” variable in RSF laws for its lack of a clear physical basis, leading to calls for models that include more explicit physical mechanisms [20, 21]. For example, the parameters that appear in eq. 1 are assumed to be constants, and they may not [22, 23, 24, 25, 26]. Furthermore, the need to regularize these laws to address numerical instabilities and ensure the physical consistency of simulations has been highlighted [27, 28]. Recently, laboratory [29] and virtual experiments [30, 31] have shown the limitations of RSF laws even in highly controlled interface conditions.
@@
Given these challenges and the limitations of existing models, there is a clear need to explore new methodologies for understanding friction dynamics. This necessity is particularly acute for modeling phenomena that span multiple scales of time and space [32]. Herein lies the potential of novel data-driven approaches, such as Recurrent Neural Networks (RNNs) [33], to offer new insights and overcome the limitations of traditional friction laws.
Neural networks have revolutionized the field of machine learning, offering powerful tools for modeling complex, non-linear relationships in data across a myriad of disciplines [34]. At their core, neural networks are composed of layers of interconnected nodes or“neurons”, each capable of performing simple computations. Through the process of training, these networks adjust their internal parameters to minimize the difference between their output and the desired outcome, effectively learning to map inputs to outputs [35].
@@
Two notable advancements in RNN architecture are the Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) networks. Both GRU and LSTM address a fundamental limitation of basic RNNs known as the vanishing gradient problem associated to recurrent connections [33], which makes it difficult for RNNs to learn dependencies between events that occur at long intervals in the input data [37, 35]. GRUs simplify the LSTM architecture while retaining its ability to capture long-term dependencies, making it both efficient and powerful for a wide range of tasks [36].
@@
This success in modeling path-dependent phenomena provides a strong foundation for the belief that similar methodologies can be applied to the study of friction. Friction, much like material plasticity, is inherently dependent on history, in its case of contact and slip between surfaces. The capability of RNNs to capture and learn from sequential, history-dependent data makes them promising candidates for advancing modeling of friction dynamics.
@@
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws based on either the aging law, eq. 2a, or the slip law eq. 2b. This approach is essentially different from the one used recently by Ishiyama and collaborators [49], who used machine learning (random forest [50]) to improve the estimation of RSF parameters eq. 1; herein, machine learning is not deployed to improve an empirical approximate law, but to replace it altogether; down the road, we envision neural-network frictional models being embedded in larger-scale simulations[21, 51, 52]. It also differs from the use made by Rouet-Leduc et al [53] and Hulbert et al [54], among others [55, 56], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. To the best of my knowledge, the closest work to ours may be the one by Wang and Sun[57], which focused on interfacial traction-opening laws, and the recent one by Cravero and K. Bhattacharya [58], who studied interface adhesion. Herein, we assume a virtual interface that is perfectly modeled by a customary aging and slip laws, eqs. 1, 2a and 2b, for which data for training and verification can be generated easily from those equations. We use those data to train a GRU network with a custom loss function to learns the dynamics without explicitly defining internal variables [47], since the state is not a physical quantity that can be measured experimentally, cf. work by Kumar and collaborators [59, 60]. In reality, the complexities of interface physics [61] are such that probably no single internal variable can be defined as long as one aims at capturing any ambitious degree of complexity. We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the hidden states of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated by the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied. Hence, when using the trained network to make predictions, one should picture an ongoing experiment that has reached a certain steady-state friction level: that information and the subsequent velocity protocol are passed to the network, which outputs a prediction for the friction evolution for the rest of the experiment.
@@
The structure of the paper is as follows: the methods section (II) introduces the GRU architecture (emphasizing how it encodes internal variable evolution), describes the generation of synthetic data for training and verification, and explains the choice of training hyperparameters, including the physics-informed loss function. The next section (III) describes the outcome of the training process, which details performance evaluation, and IV then assesses the model’s predictive accuracy on independent datasets. The discussion section (V) analyzes the findings and their implications. The final remarks section (VI) summarizes the study’s contributions and suggests future research directions.
@@ II.1 Neural networks for sequences: gated recurrent units
RNNs are engineered to handle sequential data thanks to their ability to maintain a memory of previous inputs by feeding the output of a neuron back onto itself [33]. This recursive nature allows making predictions based on the information accumulated from prior inputs. However, standard RNNs often struggle with long-term dependencies due to the vanishing gradient problem [35], where backpropagation gradients become too small or too large (over successive passings through a loop) for the network to learn effectively over many time steps [33].
Gated Recurrent Units (GRUs) are an evolution of the traditional RNN architecture, designed to mitigate the vanishing gradient problem and improve the network’s ability to learn from long sequences. GRUs introduce gated units that control the flow of information. These gates “decide” which information is relevant to keep or discard as the sequence progresses, enabling the network to maintain a longer memory, which will be useful here to handle evolution of internal variables.
The GRU mechanism is articulated mathematically through [36] the following variables:
$r_{t}$ $=\sigma(W_{ir}x_{t}+b_{ir}+W_{hr}h_{t-1}+b_{hr})$ (3a)
$z_{t}$ $=\sigma(W_{iz}x_{t}+b_{iz}+W_{hz}h_{t-1}+b_{hz})$ (3b)
$n_{t}$ $=\tanh(W_{in}x_{t}+b_{in}+r_{t}\odot(W_{hn}h_{t-1}+b_{hn}))$ (3c)
$h_{t}$ $=(1-z_{t})\odot n_{t}+z_{t}\odot h_{t-1}\,,$ (3d)
where:
  • $h_{t}$ is the hidden state at time $t$, playing a role analogous to $\theta(t)$,
  • $x_{t}$ is the input at time $t$ (in our case corresponds to $V(t)$),
  • $h_{t-1}$ is the hidden state of the layer at time $t-1$ (the loop should play the role of solving eq. 1),
  • $r_{t}$, $z_{t}$, $n_{t}$ are the reset, update, and new gates, respectively,
  • $\sigma$ is the sigmoid function, and
  • $\odot$ is the Hadamard product.
@@ II.2 Synthetic data generation
It is assumed that the state $\theta$ will not be needed by the neural network that is to replace the traditional RSF law; the internal interface evolution will be handled by $h_{t}$. This means that there will not be vectors of values of $\theta$ being used during training, but it does not imply that eqs. 2a and 2b are not being used during the datasets’ generation. Let us reiterate that we assume the hypothetical interface that is perfectly described by eq. 1, but this fact would be hidden to us if we were to analyze experimental data coming out of that interface. Having access directly to the friction coefficient is common to gouge experiments, yet not so much in other settings (e.g., PMMA friction experiments) which display more complex dynamics as stick-slip. The latter will be addressed in future work.
So, to generate the dataset for training and verifying our model, we start by defining a traditional RSF model with specific parameter values, grounded in the fundamental equations presented earlier, with particular parameter values set as follows:
@@
Using these values, we simulate 13000 random sequences to reflect diverse frictional sliding scenarios. The prior experiences with material plasticity learning [39, 44, 42] suggested that using random forcing to generate the datasets is advisable. Each sequence is designed to simulate a total sliding distance of $30D_{c}$ and is completed within 0.013 seconds. The number of velocity jumps within each sequence is randomly determined, drawing from a uniform distribution with values 2, 3, 4, 5. It is assumed that each sequence initiates from a non-zero velocity, with the initial state of the interface corresponding to the steady state for that velocity, as defined by the RSF laws: $\theta_{ss}=D_{c}/V(t=0)$.
Velocity jumps within these sequences occur at random intervals, determined by sampling times from a uniform distribution across the interval 0, 1, with 0 marking the start and 1 the end of the sliding protocol. Each sequence incorporates a hold period where the velocity drops to zero, simulating intervals where sliding stops until the subsequent jump. To avoid issues associated to dividing by zero or trying to compute the logarithm of either zero or negative numbers, a small yet finite velocity is assigned to the hold for computations ($2\cdot 10^{-9}$ m/s for aging law, $2\cdot 10^{-3}$ m/s for slip law).
@@
Having the velocity protocol, we proceed to calculate the evolution of the state variable $\theta$ using either eq. 2a or eq. 2b, and initial conditions $V(t=0)=V_{0}$ and $\theta(t=0)=\theta_{ss}$. With the velocity protocol and the evolving state variable at hand, we can then compute the friction coefficient’s evolution over time, as described by the RSF models (eq. 1).
The dataset comprises pairs of “features” and “targets”. Features represent the velocity protocol sampled at 100 equally-spaced instants during the duration of each virtual experiment, while targets correspond to variations with respect to the initial value of the friction coefficient sequence at these same instants. The reason for working with variations and not with the actual values of friction coefficient is that there is no way to predict the initial values of friction based on these datasets, they have to be provided as initial conditions over which the variations are later superimposed. There is the implicit assumption (borne by empirical observations) that the variations depend on sliding velocity, but not on the instantaneous magnitude of the friction coefficient.
For training, we normalize using the following characteristic values: $\delta\mu_{ch}=0.01$ for the friction coefficient change and $V_{ch}=0.3$ m/s for the velocities. $10\%$ of the sequences are put aside to test the network once it is trained.
When it comes to adding noise to the targets $\delta\mu_{ch}$, we choose Gaussian noise with standard deviation 0.005 and zero mean.
@@ II.3 Network details
The pyTorch python neural networks package [62] has been used.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 1 and whose parameters must be learned via training, is made up by a single GRUNet, followed by a folded Linear layer to transforms the output of the GRU through a linear transformation.